Jane ans Jill live 13.5 km apart. At 10:00 AM each girl leaves her house and jogs toward the other. The ratio of Jane's speed to Jill's is 5:4. If the girls meet at 11:15 AM, how fast does each jog? Give Rate, Time, and distance for each girl.
Answer by mananth(16946)

Jane and Jill live 13.5 km apart.
d=13.5 km
The ratio of Jane's speed to Jill's is 5:4.
let common multiple be x
Jane's speed =5x
Jill's speed =4x.
At 10:00 AM each girl leaves her house and jogs toward the other. If the girls meet at 11:15 AM,
time = 1 1/4 hour => 5/4

time = 13.5/(5x+4x)
5/4 = 13.5/9x
45x = 54
x= 6/5
5x = 5*6/5 = 6 km/h=> jane
distance = 6*1.25 = 7.50 km
4x = 4*6/5 = 4.8 km/h => Jill
distance = 1,25*4.8 =6 km





how fast does each jog? Give Rate, Time, and distance for each girl.